# Rethinking LLM Memorization through the Lens of Adversarial Compression

### v1

@@
Large language models (LLMs) trained on web-scale datasets raise substantial concerns regarding permissible data usage. One major question is whether these models ‘‘memorize’’ all their training data or they integrate many data sources in some way more akin to how a human would learn and synthesize information? The answer hinges, to a large degree, on how we define memorization. In this work, we propose the Adversarial Compression Ratio (ACR) as a metric for assessing memorization in LLMs—a given string from the training data is considered memorized if it can be elicited by a prompt shorter than the string itself. In other words, these strings can be ‘‘compressed’’ with the model by computing adversarial prompts of fewer tokens. We outline the limitations of existing notions of memorization and show how the ACR overcomes these challenges by (i) offering an adversarial view to measuring memorization, especially for monitoring unlearning and compliance; and (ii) allowing for the flexibility to measure memorization for arbitrary strings at a reasonably low compute. Our definition serves as a valuable and practical tool for determining when model owners may be violating terms around data usage, providing a potential legal tool and a critical lens through which to address such scenarios.Project page: https://locuslab.github.io/acr-memorization .
Рис. 1: We propose a compression ratio where we compare the length of the shortest prompt that elicits a training sample in response from an LLM. If a string in the training data can be compressed, i.e. the minimal prompt is shorter than the sample, then we call it memorized. Our test is an easy to describe tool that is useful in the effort to gauge the misuse of data.
@@ 1 Introduction
A central question in the discussion of large language models (LLMs) concerns the extent to which they memorize their training data versus the ways in which they generalize to new tasks and settings. Most practitioners seem to (at least informally) believe that LLMs do some degree of both: they clearly memorize some aspects of the training data—for example, are often able to reproduce large portions of training data verbatim (Carlini et al., 2023)—but they also seem to learn from this data allowing them to generalize to new settings. The precise extent to which they do one or the other has massive implications for the practical and legal aspects of such models (Cooper et al., 2023). Do LLMs truly produce new content, or do they only remix their training data? Should the act of training on copyrighted data be deemed unfair use of data, or should fair use be judged by the model’s memorization? With regard to people, we distinguish plagiarising content from learning from it, but how should this extend to LLMs? The answer to such questions inherently relates to the extent to which LLMs memorize their training data.
@@
However, even defining memorization for LLMs is challenging and many existing definitions leave a lot to be desired. Certain formulations claim that a passage from the training data is memorized if the LLM can reproduce it exactly (Nasr et al., 2023). However, this ignores situations where, for instance, a prompt instructs the model to exactly repeat some phrase. Other formulations define memorization by whether or not prompting an LLM with a portion of text from the training set results in the completion of that training datum (Carlini et al., 2023). But these formalisms rely fundamentally on the completions being a certain size, and typically very lengthy generations are required for sufficient certainty of memorization. More crucially, these definitions are too permissive because they ignore situations where model developers can (for legal compliance) post-hoc ‘‘align’’ an LLM by instructing their models not to produce certain copyrighted content (Ippolito et al., 2023). But has such an instructed model really not memorized the sample in question, or does the model still contain all the information about the datum in its weights and hides behind an illusion of compliance? Asking such questions becomes critical because this illusion of ‘‘unlearning’’ can often be easily broken as we show in Sections 4.1 and 4.3.
@@
In this work, we propose a new definition of memorization based on a compression argument. Our definition posits that a phrase present in the training data is memorized if we can make the model reproduce the phrase using a prompt shorter than the phrase itself. Operationalizing this definition requires finding the shortest adversarial input prompt that is specifically optimized to produce a target output. We call this ratio of input to output tokens the Adversarial Compression Ratio (ACR). In other words, memorization is inherently tied to whether a certain output can be represented in a compressed form, beyond what language models can do with typical text. We argue that such a definition provides an intuitive notion of memorization—if a certain phrase exists within the LLM training data (e.g., is not itself generated text) and it can be reproduced with fewer input tokens than output tokens, then the phrase must be stored somehow within the weights of the LLM. Although it may be more natural to consider compression in terms of the LLM-based notions of input/output perplexity, we argue that a simple compression ratio based on input/output token counts provides a more intuitive explanation to non-technical audiences, and has the potential to serve as a legal basis for important questions about memorization and permissible data use.
@@
In addition to its intuitive nature, our definition has several other desirable qualities. Experimentally, we show that it appropriately ascribes many famous quotes as being memorized by existing LLMs, that is, their ACR is greater than one. On the other hand, we find that text not in the training data of an LLM, such as samples posted on the internet after the training period, are not compressible, that is their ACR is less than one.
@@
We examine several so-called unlearning methods through this lens to show that they do not substantially affect the memorization of the underlying model. That is, even after explicit instruction- or fine-tuning, models asked to ‘‘forget’’ certain pieces of content are still able to reproduce them with a high ACR— in fact, not much smaller than with the original model. Our approach provides a simple and practical perspective on what memorization can mean, providing a useful tool for functional and legal analysis of LLMs.
@@ 2 Do We Really Need Another Notion of Memorization?
With LLMs ingesting more and more data, questions about their memorization are attracting attention (e.g. Carlini et al., 2019; 2023; Nasr et al., 2023; Zhang et al., 2023). There remains a pressing need to accurately define memorization in a way that serves as a practical tool to ascertain the fair use of public data from a legal standpoint. To ground the problem, consider the court’s role in determining whether an LLM is breaching copyright. What constitutes a breach of copyright remains contentious and prior work defines this on a spectrum from ‘training on a data point itself constitutes violation’ to ‘copyright violation only occurs if you verbatim regurgitate training data’. To formalize our arguments and more carefully frame our points, we use three definitions as a starting point for our summary of the field thus far.
Рис. 2: Adversarial Compression is harder to hack, better attuned to what problematic data usage is, and requires no parameter choice. This makes it a better choice than alternative tests for fair use. If we deploy tests that look for exact auto-completion, we will flag very few cases as problematic. On the other hand, if we say any training done on copyrighted data is an issue we will have a very restrictive system. We argue for defining fair use with adversarial compression.
Definition 1 (Discoverable Memorization (Carlini et al., 2023))
Given a generative model $M$, a sample $y$ made of a prefix $y_{\text{prefix}}$ and a suffix $y_{\text{suffix}}$ from the training data is discoverably memorized if the prefix elicits the suffix in response, or $M(y_{\text{prefix}})=y_{\text{suffix}}$.
@@
    Very Permissive: This requires a model to output an exact match of some number of tokens from the training set (under greedy decoding). This is extremely permissive and misses cases when the exact substrings are the second most likely choice.
@@
    Sensitive to Parameter Choice: We need to choose several words (or tokens) to include in the prompt and a number of tokens that have to match exactly in the output to turn this definition into a practical binary test for memorization. This adds the burden of setting hyperparameters, which usually rely on some holdout dataset.
Definition 2 (Extractable Memorization (Nasr et al., 2023))
Given a generative model $M$, a sample $y$ from the training data is extractably memorized if an adversary without access to the training set can find an input prompt $p$ that elicits $y$ in response, or $M(p)=y$.
The next definition says a string is Extractably Memorized if there exists a prompt that elicits the string in response. This falls too far on the other side of the issue by being very restrictive—what if the prompt includes the entire string in question, or worse the instructions to repeat it? LLMs that are good at repeating will follow that instruction and output any string they are asked to. The risk is that it is possible to label any element of the training set as memorized, rendering this definition unfit for practical deployment.
Definition 3 (Counterfactual Memorization (Zhang et al., 2023))
Given a training algorithm $A$ that maps a dataset $\mathcal{D}$ to a generative model $M$ and a measure of model performance $S(M,y)$ on a specific sample $y$, the counterfactual memorization of a training example $y\in\mathcal{D}$ is given by:
$\texttt{mem}(y):=\underbrace{\mathbb{E}_{D\subset\mathcal{D},y\in D}[S(A(D),y)]}_{\text{performance on $y$ when trained with $y$}}-\underbrace{\mathbb{E}_{D^{\prime}\subset\mathcal{D},y\not\in D^{\prime}}[S(A(D^{\prime}),y)]}_{\text{performance on $y$ when {not} trained with $y$}},$
@@
The third notion of memorization aims to separate memorization from generalization and is called Counterfactual Memorization, which requires a test that includes retraining many models. Given, the cost of retaining large language models, such a definition remains impractical for legal use.
@@
Membership Inference Attacks (MIA) are designed to determine whether a data point is in the training set of a given model, which is a common focus of privacy research (e.g. Shokri et al., 2017). Upon first glance, MIAs may look like tests for memorization and they are even intimately related to auditing machine unlearning (Carlini et al., 2021; Pawelczyk et al., 2023). However, there is a subtle and crucial difference between membership and memorization. In particular, the ongoing lawsuits in the field (e.g. as covered by Metz & Robertson, ) leave open the possibility that reproducing another’s creative work is problematic but training on samples from that data may not be. This is common practice in the arts—consider that a copycat comedian telling someone else’s jokes is stealing, but an up-and-comer learning from tapes of the greats is doing nothing wrong. In particular, this approach also has three issues.
@@
    Very Restrictive: LLMs are typically trained on trillions of tokens. Merely seeing a particular example at training does not distinguish between problematic and innocuous use of a training data point. Akin to plagiarism, it is okay to read copyrighted books, it is only problematic when content is copied. We need a similarly permissive definition of memorization for LLMs.
@@
    Hard to Arbitrate: Arbitrating the validity of an MIA is itself problematic because it assumes good faith from the side of a corporation in releasing information about the data that they trained on in front of an arbiter. This becomes problematic given the inherently adversarial nature of such an event, and the potential of plausible deniability.
@@
    Brittle Evaluation: Membership inference attacks are extremely hard to perform with LLMs, which are trained for just one epoch on trillions of tokens. Some recent work shows that LLM membership inference is extremely brittle (Duan et al., 2024), which is supported by prior work theoretically examining the success of MIAs as dataset sizes increase (Maini et al., 2021).
Perplexity is complex and hackable
Another notion that appeals to the information theorist is the use of perplexity. We omit a formal definition here for brevity, but this encompasses approaches that use the model as a probability distribution over tokens to compute the information content of a string and estimate its optimal compression rate under the given model. Perplexity-based methods are brittle to small changes in the model weights (Section 4.1) and lack the approachability of our definition. It is easier to picture a non-technical governing body and the broader community understanding and using a test based on input-output compression than one based on advanced information theory.
@@ 3 How to Measure Memorization with Adversarial Compression
Our definition of memorization is based on answering the following question: Given a piece of text, is the minimal prompt that elicits that text exactly shorter than the sample itself? In this section, we lay the definition formally and we introduce our MiniPrompt algorithm that we use to answer our central question.
@@ 3.1 A New Definition of Memorization
To begin, let a target natural text string $s$ have a token sequence representation $x\in\mathcal{V}$ which is a list of integer-valued indices that index a given vocabulary $\mathcal{V}$. We use ||$\cdot$|| ⋅ | to count the length of a token sequence. A tokenizer $T:s\mapsto x$ maps from strings to token sequences. Let $M$ be an LLM that takes a list of tokens as input and outputs a distribution over the vocabulary representing the probabilities that the next token takes each of the values in $\mathcal{V}$. Consider that $M$ can perform generation by repeatedly predicting the next token from all the previous tokens with the argmax of its output appended to the sequence at each step (this process is called greedy decoding). With a slight abuse of notation, we will also call the greedy decoding result the output of M. Let $y$ be the token sequence generated by $M$, which we call a completion or response:
$y=M(x).$ (1)
To describe Equation 1 in natural language we say that the model generates $y$ when prompted with $x$ or that $x$ elicits $y$ as a response from $M$. Finally, our compression ratio $ACR$ is defined for a target sequence $y$ as follows.
$ACR=\frac{|y|}{|x^{*}|},\text{ where, }x^{*}=\operatorname*{arg\,min}_{x}|x|\text{ s.t. }M(x)=y.$ (2)
Definition 4 (Compressible Memorization)
@@
Given a generative model $M$, a sample $y$ from the training data is memorized if an adversary can find an input prompt $x$ of fewer tokens than $y$ that elicits $y$ in response, or $M(x)=y$ and $|x|<|y|$.
@@
Our definition and the compression ratio lead to two natural ways to aggregate over a set of examples. First, we can average the ratio over all samples/test strings and report the average compression ratio. Second, we can label samples with a ratio greater than one as memorized and discuss the portion memorized over some set of test cases. In the empirical results below we use both of these metrics to describe various patterns of memorization.
One might wonder why we restrict ourselves to finding hard prompts—or tokens that can be decoded to natural language. Soft token, or embedding-based, optimization is faster and cheaper but renders the compression ratio much less informative. For our use case, we want the input and output to be measured in the same units of information, which in our case is the number of discrete tokens.
3.2 MiniPrompt: A Practical Test for Compressible Memorization

### v2

@@
Large language models (LLMs) trained on web-scale datasets raise substantial concerns regarding permissible data usage. One major question is whether these models ‘‘memorize’’ all their training data or they integrate many data sources in some way more akin to how a human would learn and synthesize information. The answer hinges, to a large degree, on how we define memorization. In this work, we propose the Adversarial Compression Ratio (ACR) as a metric for assessing memorization in LLMs. A given string from the training data is considered memorized if it can be elicited by a prompt (much) shorter than the string itself—in other words, if these strings can be ‘‘compressed’’ with the model by computing adversarial prompts of fewer tokens. The ACR overcomes the limitations of existing notions of memorization by (i) offering an adversarial view of measuring memorization, especially for monitoring unlearning and compliance; and (ii) allowing for the flexibility to measure memorization for arbitrary strings at a reasonably low compute. Our definition serves as a practical tool for determining when model owners may be violating terms around data usage, providing a potential legal tool and a critical lens through which to address such scenarios. Project page: https://locuslab.github.io/acr-memorization
Рис. 1: We propose a compression ratio where we compare the length of the shortest prompt that elicits a training sample in response from an LLM to the length of that sample. If a string in the training data can be compressed, i.e. the minimal prompt is shorter than the sample, then we call it memorized. Our test is an easy-to-describe tool that is useful in the effort to gauge the misuse of data.
@@ 1 Introduction
A central question in the discussion of large language models (LLMs) concerns the extent to which they memorize their training data versus how they generalize to new tasks and settings. Most practitioners seem to (at least informally) believe that LLMs do some degree of both: they clearly memorize parts of the training data—for example, are often able to reproduce large portions of training data verbatim (Carlini et al., 2023)—but they also seem to learn from this data, allowing them to generalize to new settings. The precise extent to which they do one or the other has massive implications for the practical and legal aspects of such models (Cooper et al., 2023). Do LLMs truly produce new content, or do they only remix their training data? Should the act of training on copyrighted data be deemed unfair use of data, or should fair use be judged by the model’s memorization? With respect to people, we distinguish plagiarising content from learning from it, but how should this extend to LLMs? The answer to such questions inherently relates to the extent to which LLMs memorize their training data.
@@
However, even defining memorization for LLMs is challenging and many existing definitions leave a lot to be desired. Certain formulations claim that a passage from the training data is memorized if the LLM can reproduce it exactly (Nasr et al., 2023). However, this ignores situations where, for instance, a prompt instructs the model to exactly repeat some phrase. Other formulations define memorization by whether or not prompting an LLM with a portion of text from the training set results in the completion of that training datum (Carlini et al., 2023). But these formalisms rely fundamentally on the completions being a certain size, and typically very lengthy generations are required for sufficient certainty of memorization. More crucially, these definitions are too permissive because they ignore situations where model developers can (for legal compliance) post-hoc ‘‘align’’ an LLM by instructing their models not to produce certain copyrighted content (Ippolito et al., 2023). But has such an instructed model really not memorized the sample in question, or does the model still contain all the information about the datum in its weights while it hides behind an illusion of compliance? Asking such questions becomes critical because this illusion of ‘‘unlearning’’ can often be easily broken as we show in Sections 4.1 and 4.3.
@@
In this work, we propose a new definition of memorization based on a compression argument. Our definition posits that a phrase present in the training data is memorized if we can make the model reproduce the phrase using a prompt (much) shorter than the phrase itself. Operationalizing this definition requires finding the shortest adversarial input prompt that is specifically optimized to produce a target output. We call this ratio of input to output tokens the Adversarial Compression Ratio (ACR). In other words, memorization is inherently tied to whether a certain output can be represented in a compressed form, beyond what language models can do with typical text. We argue that such a definition provides an intuitive notion of memorization—if a certain phrase exists within the LLM training data (e.g., is not itself generated text) and it can be reproduced with fewer input tokens than output tokens, then the phrase must be stored somehow within the weights of the LLM. Although it may be more natural to consider compression in terms of the LLM-based notions of input/output perplexity, we argue that a simple compression ratio based on input/output token counts provides a more intuitive explanation to non-technical audiences, and has the potential to serve as a legal basis for important questions about memorization and permissible data use.
@@
In addition to its intuitive nature, our definition has several other desirable qualities. We show that it appropriately ascribes many famous quotes as being memorized by existing LLMs (i.e. they have high ACR values). On the other hand, we find that text not in the training data of an LLM, such as samples posted on the internet after the training period, are not compressible, that is their ACR is low.
@@
We examine several unlearning methods using ACR to show that they do not substantially affect the memorization of the model. That is, even after explicit finetuning, models asked to ‘‘forget’’ certain pieces of content are still able to reproduce them with a high ACR—in fact, not much smaller than with the original model. Our approach provides a simple and practical perspective on what memorization can mean, providing a useful tool for functional and legal analysis of LLMs.
@@ 2 Do We Really Need Another Notion of Memorization?
With LLMs ingesting more and more data, questions about their memorization are attracting attention (e.g. Carlini et al., 2019, 2023; Nasr et al., 2023; Zhang et al., 2023). There remains a pressing need to accurately define memorization in a way that serves as a practical tool to ascertain the fair use of public data from a legal standpoint. To ground the problem, consider the court’s role in determining whether an LLM is breaching copyright. What constitutes a breach of copyright remains contentious and prior work defines this on a spectrum from ‘training on a data point itself constitutes violation’ to ‘copyright violation only occurs if a model verbatim regurgitates training data’. To formalize our argument for a new notion of memorization, we start with three definitions from prior work to highlight some of the gaps in the current thinking about memorization.
Definition 1 (Discoverable Memorization (Carlini et al., 2023)).
Given a generative model $M$, a sample $y$ from the training data made of a prefix $y_{\text{prefix}}$ and a suffix $y_{\text{suffix}}$ is discoverably memorized if the prefix elicits the suffix in response, or $M(y_{\text{prefix}})=y_{\text{suffix}}$.
@@
    Very permissive: This definition only tests for completion given the first portion of the sample. Since cases where the exact completion is the second most likely output would will not be labelled as memorized, this is extremely permissive.
@@
    Parameter choice requires validation data: We need to choose several words (or tokens) to include in the prompt and a number of tokens that have to match exactly in the output to turn this definition into a practical binary test for memorization. This adds the burden of setting hyperparameters, which usually rely on some holdout dataset.
Definition 2 (Extractable Memorization (Nasr et al., 2023)).
Given a generative model $M$, a sample $y$ from the training data is extractably memorized if an adversary, without access to the training set, can find an input prompt $p$ that elicits $y$ in response, or $M(p)=y$.
A string is extractably memorized if there exists a prompt that elicits the string in response. This falls too far on the other side of the issue by being very restrictive—what if the prompt includes the entire string in question, or worse, the instructions to repeat it? LLMs that are good at repeating will follow that instruction and output any string they are asked to. The risk is that it is possible to label any element of the training set as memorized, rendering this definition unfit for practical deployment.
Definition 3 (Counterfactual Memorization (Zhang et al., 2023)).
Given a training algorithm $A$ that maps a dataset $D\subset\mathcal{D}$ to a generative model $M$ and a measure of model performance $S(M,y)$ on a specific sample $y$, the counterfactual memorization of a training example $y\in\mathcal{D}$ is given by:
$\texttt{mem}(y):=\underbracket{\mathbb{E}_{D\subset\mathcal{D},y\in D}[S(A(D),y)]}_{\text{performance on y when trained with y}}-\underbracket{\mathbb{E}_{D^{\prime}\subset\mathcal{D},y\not\in D^{\prime}}[S(A(D^{\prime}),y)]}_{\text{performance on y when {not} trained with y}},$
@@
Counterfactual memorization aims to separate memorization from generalization and requires a test that includes retraining many models. Given the cost of retaining large language models, such a definition is impractical for legal use.
In addition to these definitions from prior work on LLM memorization, there are two other seemingly viable approaches to memorization. Ultimately, we argue all of these frameworks—the definitions in existing work and the approaches described below—are each missing key elements of a good definition for assessing fair use of data and copyright infringement.
@@
Perhaps if a copyrighted piece of data is in the training set at all we might consider it a problem. However, there is a subtle but crucial difference between training set membership and memorization. In particular, the ongoing lawsuits in the field (e.g. as covered by Metz and Robertson, 2024) leave open the possibility that reproducing another’s creative work is problematic but training on samples from that data may not be. This is common practice in the arts—consider that a copycat comedian telling someone else’s jokes is stealing, but an up-and-comer learning from tapes of the greats is doing nothing wrong. So while membership inference attacks (MIAs) (e.g. Shokri et al., 2017) may look like tests for memorization and they are even intimately related to auditing machine unlearning (Carlini et al., 2021; Pawelczyk et al., 2023; Choi et al., 2024), they have three issues as tests for memorization.
@@
    Very restrictive: LLMs are typically trained on trillions of tokens. Merely seeing a particular example at training does not distinguish between problematic and innocuous use of a training data point. Akin to plagiarism, it is okay to read copyrighted books, but copying is problematic.
@@
    Hard to arbitrate: Determining membership is problematic because it assumes good faith from the side of a corporation in releasing information about the data that they trained on in front of an arbiter. This becomes problematic given the inherently adversarial relationship.
@@
    Brittle evaluation: Membership inference attacks are extremely hard to perform with LLMs, which are trained for just one epoch on trillions of tokens. Some recent work shows that LLM membership inference is extremely brittle (Duan et al., 2024; Maini et al., 2021).
Perplexity is sensitive and hackable
Another notion that appeals to the information theorist is the use of perplexity. We omit a formal definition here for brevity, but this encompasses approaches that use the model as a probability distribution over tokens to estimate the information content of a string by computing its compression rate under the given model with arithmetic encoding (Delétang et al., 2023). Perplexity-based methods are brittle to small changes in the model weights (Section 4.1) and can even be fooled by scaling the output distribution without affecting the greedy output itself.
@@ 3 How to Measure Memorization with Adversarial Compression
Our definition of memorization is based on answering the following question: Given a piece of text, how short is the minimal prompt that elicits that text exactly? In this section, we formally define and introduce our MiniPrompt algorithm that we use to answer our central question.
@@ 3.1 A New Definition of Memorization
To begin, let a target natural text string $s$ have a token sequence representation $x\in\mathcal{V}^{*}$ which is a list of integer-valued indices that index a given vocabulary $\mathcal{V}$. We use ||$\cdot$|| ⋅ | to count the length of a token sequence. A tokenizer $T:s\mapsto x$ maps from strings to token sequences. Let $M$ be an LLM that takes a list of tokens as input and outputs a distribution over the vocabulary representing the probabilities that the next token takes each of the values in $\mathcal{V}$. Consider that $M$ can perform generation by repeatedly predicting the next token from all the previous tokens with the argmax of its output appended to the sequence at each step (this process is called greedy decoding). With a slight abuse of notation, we will also call the greedy decoding result the output of $M$. Let $y$ be the token sequence generated by $M$, which we call a completion or response: $y=M(x)$, which in natural language says that the model generates $y$ when prompted with $x$ or that $x$ elicits $y$ as a response from $M$. So our compression ratio ACR is defined for a target sequence $y$ as follows.
$\textsc{ACR}{}(M,y)=\frac{|y|}{|x^{*}|},\text{ where, }x^{*}=\operatorname*{arg\,min}_{x}|x|\text{ s.t. }M(x)=y.$ (1)
Definition 4 ($\tau$-Compressible Memorization).
@@
Given a generative model $M$, a sample $y$ from the training data is $\tau$-memorized if the $\textsc{ACR}{}(M,y)>\tau(y)$.
The threshold $\tau(y)$ is a configurable parameter of this definition. We might choose to compare the ACR to the compression ratio of the text when run through a general-purpose compression program (explicitly assumed not to have memorized any such text) such as GZIP (Gailly and Adler, 1992) or SMAZ (Sanfilippo, 2006). This amounts to setting $\tau(y)$ equal to the SMAZ compression ratio of $y$, for example. Alternatively, one might even use the compression ratio of the arithmetic encoding under another LLM as a comparison point, for example if it was known with certainty that the LLM was never trained on the target output, and hence could not have memorized it (Delétang et al., 2023). In reality, copyright attribution cases are always subjective, and the goal of this work is not to argue for the right threshold function, rather to advocate for the adversarial compression framework for arbitrating fair data use. Thus, we use $\tau=1$ in the experiments below, which we believe has substantial practical value. [^1] There exist prompts like “count from $1$ to $1000$,” for which a chat model $M$ is able to generate “$1,2,\ldots,1000$,” which results in a very high ACR. However, for copyright purposes, we argue that this category of algorithmic prompts are in the gray area where determining memorization is difficult and beyond the scope of this paper given our primary application to creative works.
@@
Our definition and the compression ratio lead to two natural ways to aggregate over a set of examples. First, we can average the ratio over all samples/test strings and report the average compression ratio (this is $\tau$-independent). Second, we can label samples with a ratio greater than one as memorized and discuss the portion memorized over some set of test cases (for our choice of $\tau=1$). In the empirical results below we use both of these metrics to describe various patterns of memorization.
3.2 MiniPrompt: A Practical Algorithm for Compressible Memorization